Hydrofluoric acid in produced by a reaction between hydrogen and fluorine gases. Given 23 g of hydrogen to start, what volume of hydrofluoric acid Will ve produced at standard pressure and temperature of 28 degrees Celsius. Assume excess fluorine gas.
1
Expert's answer
2019-04-09T07:54:41-0400

Hydrofluoric acid is produced in the following reaction:

H2 + F2 = 2HF

Molecular weight of H2 = 2 g/mol

In the reaction, 23 g of hydrogen (H2) is used. As a result, 23 g / 2 g/mol = 11.5 moles of hydrogen are used to produce 11.5 / 2 = 5.75 moles of HF.

As at 28°C, hydrogen fluoride is a gas, the ideal gas equation should be used to estimate the volume of the gas.

According to the ideal gas equation:

V = nRT / P,

where V - gas volume, n - number of moles, R - gas constant, T - temperature, P - pressure.

Here, n = 5.75 moles, R = 0.082057 L•atm•K-1•mol-1, T = 28°C = 301.15 K, P = 1 atm.

As a result,

V = 5.75 moles × 0.082057 L•atm•K-1•mol-1× 301.15 K / 1 atm= 142 L.

Answer: 142 L.